OpenStudy (anonymous):

help, pretty please?(: Which ordered pairs in the form (x, y) are solutions to the equation (x + 3)y = 14? Choose all answers that are correct. A. (−5, −7) B. (−1, 7) C. (10, −2) D. (11, 1)

OpenStudy (anonymous):

lets check the first one

OpenStudy (anonymous):

is \[(-5+3)(-7)=14\]?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

its A, b & D... drop the #'s in and calculate they all come out to 14

OpenStudy (anonymous):

actually "yes" \[(-5+3)\times (-7)=-2\times -7=14\]

OpenStudy (anonymous):

thank you! @khusker (: